OpenStudy (anonymous):

If you are making 28 brownies and the recipe calls for 2/3 cup of butter for a serving of 16 brownies how much butter will you use? (type the integer, fraction, or mixed number)

OpenStudy (kkutie7):

\[28brownies*\frac{\frac{2}{3}cup of butter}{16brownies}=?\]

OpenStudy (kkutie7):

so if you look it at it this way your units (brownies) are being cancelled leaving you with cups as the only unit. you calculate it like this: \[\frac{28*\frac{2}{3}}{16}=cups of butter\]

OpenStudy (triciaal):